# An introduction to reinforcement learning for neuroscience

### v2

@@ 4 Q-learning
In some cases, we may not know the transition function or it could be expensive to simulate. Additionally, research suggests that dopamine activity reflects not just state values but also action values (Roesch et al., , 2007; Morris et al., , 2006). This suggests an alternative algorithm for biological learning, where animals learn state-action values in what is known as ‘Q-learning’. Such Q-learning models have been commonly used to explain animal behaviour and neural activity (Niv, , 2009; Mattar and Daw, , 2018).
To understand this algorithm, we first need to define the state-action values under policy $\pi$, $Q^{\pi}(s,a)$:
$Q^{\pi}(s,a)=\mathbb{E}_{\tau\sim p_{\pi}(\tau)}\left[\sum_{t^{\prime}=t}\gamma^{t^{\prime}-t}r_{t^{\prime}}|s_{t}=s,a_{t}=a\right].$ (12)
Given $Q^{\pi}$, it is simple to perform action selection by taking the action with the highest expected reward in a given state,
$a^{*}(s)=\text{argmax}_{a}Q^{\pi}(s,a).$ (13)
To learn the $Q$-values necessary for such action selection, we start by expanding Equation 12,
@@
Importantly, this self-consistency expression only holds when the Q-values have converged to the true expected rewards, and the associated greedy policy is therefore optimal (Sutton and Barto, , 2018). We can now use Equation 15 as an objective by defining
$\mathcal{L}\propto\left(Q(s,a)-(r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a)}\left[\text{max}_{a^{\prime}}Q(s^{\prime},a^{\prime})\right])\right)^{2},$ (16)
@@
This is the so-called Q-learning update rule (Watkins, , 1989; Figure 3A), where we have estimated the expectation with the single sample actually seen by the agent in the last line.
@@
Q-learning is guaranteed to converge to the optimal policy in the limit of infinitesimal learning rates and infinite sampling of the state-action space (Watkins and Dayan, , 1992; Sutton and Barto, , 2018). However, following the greedy policy $a^{*}(s)=\text{argmax}_{a}Q(s,a)$ before convergence of the Q-values can lead to undersampling of the state-action space and poor performance. It is therefore common to either use an ‘$\epsilon$-greedy’ policy, $\pi(a|s)=\epsilon/|\mathcal{A}|+(1-\epsilon)I_{a=a^{*}(s)}$, or a softmax-policy, $\pi(a|s)\propto\text{exp}(\beta Q(s,a))$, to collect the experience used to update the Q-values (Figure 3B). Such exploration strategies and their biological correlates are discussed in more detail in Section 9.
These approaches make Q-learning an ‘off-policy’ algorithms, since the policy used in the learning update (the greedy policy) is different from the policy used for action selection (the stochastic policy). An on-policy version of Q-learning known as ‘SARSA’ (state-action-reward-state-action) is also commonly used, where the Q-learning update uses the Q-value corresponding to the action $a_{t+1}$ sampled at the next timestep instead of the greedy action (Figure 3C):
@@
This will converge to the true Q-values for a given policy $\pi$, similar to how the TD learning rule in Equation 9 is guaranteed to converge to the true value function for a given policy, again under assumptions of infinite sampling of the space.
In settings where animals have to choose between multiple actions with different values, studies have found evidence for midbrain dopamine neurons encoding either quantities resembling the prediction error used for Q-learning (Roesch et al., , 2007; Niv, , 2009) or SARSA (Morris et al., , 2006; Niv, , 2009). These findings suggest that they may not just be abstract learning algorithms, but instead have pluasible implementations in biological neural circuits. However, the methods considered so far also have notable shortcomings. For example, the amount of data needed to learn state(-action) values and the assumption of a stationary environment can be prohibitive for animals needing to act in a rapidly changing world, where bad decisions have fatal consequences.
@@ 5 Model-free and model-based reinforcement learning
We have so far considered what is known as ‘model-free’ reinforcement learning algorithms. These involve learning a stimulus-response pattern that says ‘when in state $s$, take action $a$’. Such algorithms do not require much computation at decision time, where they rely on cached state or action values. However, it can require a lot of experience with the environment to learn these model-free policies, and they can be inflexible in changing environments. This is incompatible with many aspect of animal behaviour, which we know is adaptive and can benefit from ‘latent learning’ in an environment before a reward-driven task is ever encountered (Blodgett, , 1929; Tolman, , 1948).
On the other hand, ‘model-based’ reinforcement learning uses a model of the world to simulate the consequences of different actions at decision time. This can be much more data efficient, since learning a world model is often simpler than learning a full policy (Figure 4A). In machine learning settings, model-based RL has exhibited impressive performance across a range of tasks with large state spaces, including Atari, chess, shogi, and Go (Silver et al., , 2018; Schrittwieser et al., , 2020; Deisenroth and Rasmussen, , 2011). In a biological context, the idea of first learning a model of the environment, and then using it to guide reward-driven behaviour also provides one plausible explanation for latent learning and other types of rapid adaptation. However, model-based decision making can be computationally intensive at decision-time, which is a challenge for animals that rely on rapid decision making for survival (Figure 4B).
@@
If the environment is determinstic, $p(s^{\prime}|s,a)$ is a delta function, and otherwise the next-state expectations may need to be approximated with multiple samples. Unfortunately, optimizing over all possible action sequences in Equation 20 is in general an exponentially large search problem in the planning depth, which makes it infeasible for any reasonably sized problem. It is therefore common to either use ‘depth-first search’ with limited breadth, or ‘breadth-first search’ with limited depth. In breadth-first search, we consider all possible actions at each level of the search tree but terminate the search at a finite depth, instead using cached ‘model-free’ state-values to estimate the reward-to-go from the terminal states. Such ‘plan-until-habit’ has also been proposed as a model of human behaviour (Keramati et al., , 2016). In depth-first search, we instead sample a series of paths from $s_{t}$ to termination (or some upper bound), using a heuristic to prioritize actions expected to be good, and then pick an action with high expected reward (Huys et al., , 2012).
@@
For both of these strategies, it is necessary to trade off the temporal opportunity cost of planning with the improvement in policy and expected reward (Botvinick and Cohen, , 2014; Agrawal et al., , 2022). This has been a popular research area in cognitive science, where a wealth of literature on ‘resource-rational’ decision making has emerged in recent years (Griffiths et al., , 2019; Callaway et al., , 2022). However, this literature has often considered optimal behaviour in simple tasks with less focus on the learning process and neural mechanisms that might implement the necessary computations. Bridging this gap, recent work has suggested that frontal cortex and striatum might initially store a ‘model-free’ policy in its network state, which is gradually updated with model-based information from the hippocampal formation until the policy improvement is outweighed by the temporal opportunity cost of planning (Jensen et al., , 2023).
@@
While several model-based and model-free reinforcement learning methods have thus been developed and used to model animal learning and behaviour, it remains an open question when and whether these different strategies drive animal behaviour. A different line of research has therefore explicitly investigated the balance between model-based and model-free RL in biological agents (Daw et al., , 2005; Geerts et al., , 2020; Lengyel and Dayan, , 2007), where the choice between the two approaches is thought to be guided by some notion of optimality on the basis of available resources and uncertainty about the environment. A popular paradigm for these studies is the so-called ‘two-step’ task developed by Daw and colleagues (Daw et al., , 2011; Momennejad et al., , 2017; Wang et al., , 2018; although note Akam et al., , 2015).
@@
Such work has shown that animals can use both model-free and model-based decision making, with the dorsolateral striatum being particularly important for model-free reinforcement learning (Yin et al., , 2004, 2005), and the dorsomedial striatum, prefrontal cortex, and hippocampal formation being important for model-based decision making (Vikbladh et al., , 2019; Geerts et al., , 2020; Miller et al., , 2017; Niv, , 2009; Killcross and Coutureau, , 2003). This also has interesting parallels to recent work in motor learning, where the basal ganglia were found to be sufficient for ‘habitual’ motor sequences even in the absence of motor cortex, while motor cortex was necessary for more flexible motor behaviours that are likely to require a high-level ‘schema’ of the task structure (Mizes et al., 2023b, ; Mizes et al., 2023a, ). In Section 9 we will see how combining these model-based and model-free ideas with deep learning can lead to human-level performance in tasks such as chess and Go that require long-term planning.
@@ 6 The successor representation
As we saw in the previous section, an important distinction can be made between model-free reinforcement learning methods, which cache stimulus-response mappings based on prior experience, and model-based reinforcement learning methods, which compute a policy by simulating possible futures using a world model at decision-time. However, we have also noted how animals both need the flexibility of model-based methods as well as the rapid decision making afforded by model-free methods. It has therefore been suggested that animals use intermediate methods that combines some model-free and some model-based features. A particularly prevalent theory has been that of the ‘successor representation’ (SR), which has been proposed to explain both human behaviour (Momennejad et al., , 2017) and features of neural activity (Stachenfeld et al., , 2017). In particular, the SR allows for flexible adaptation to changing reward functions without having to perform expensive simulations at decision time.
@@
Figure 4: Model-based reinforcement learning. (A) Learning curves for model-based (MB) and model-free (MF) RL agents. The MB agent used depth-first search to compute an optimal path at each decision point, gradually learning the reward and transition functions while exploring the environment. The MF agent was a simple Q-learning agent with $\epsilon=0$ and learning rate $\alpha=1$. (B) Wallclock time needed to run 100 episodes of cliffworld with either the MB or MF agents from (A), as a function of the length of the environment. While the MB agent required less experience to learn a good policy, the wallclock time per episode was much larger than for the MF agent. This illustrates an important balance between model-based and model-free reinforcement learning, where MF methods usually require more experience but MB methods require more compute at decision time. (C) Learning curve for an agent using the successor representation (SR) together with learning curves for the model-based agent in (A) and the greedy TD-agent from Figure 2. The goal was moved from location (9, 0) to location (0, 4) at episode 40 (vertical black line), and location (9, 0) was instead given a reward of -5. The MB and SR agents had their reward functions updated to reflect this change and rapidly adapted their policies, while the TD agent had no such mechanism for robustness to changing reward functions. Reward curves were convolved with a Gaussian kernel ($\sigma=3$ episodes), which is why performance appears to decrease slightly before episode 40. The TD and SR agents were assumed to have access to a 1-step world model at initialization, while the MB agent learned the transition structure from experience. (D) SR agents cannot always adapt to new reward functions if the newly rewarded states have low probablity under the old policy. Left column: Value functions are shown for an agent that learned an initial policy in an environment with a small reward in the upper left corner and intermediate reward in the upper right corner. The agent learned to make an initial rightward choice in this environment (grey arrows). Right column: A new large reward was then introduced in either the top left corner (top row) or bottom left corner (bottom row) and the value function recomputed (Equation 29). In the second case, the agent was unable to adapt since the old policy had low probability of reaching the large reward, even after initially going to the left. This results in a low expected value for going left from the start state (red circle), and a suboptimal policy that continues to go right (red arrow). (E) Learning curve for a standard Q-learning agent (blue) or DYNA agents that perform different numbers of Q-value updates after each physical action (legend). These DYNA updates used cached experience rather than data from a learned world model. DYNA agents make better use of limited experience at the cost of increased compute (proportional to the number of updates).
The SR (Dayan, , 1993) rewrites the expected reward starting from state $s$ as:
@@
This provides a better starting point than the old policy and reward function (Figure 4C), but the SR can sometimes fail to generalize since the new value function will eventually lead to a change in policy and therefore having to update $\bm{M}$ (Figure 4D). The successor matrix can be learned by simple temporal-difference learning when transitioning from $s_{t}$, analogous to Equation 9:
@@
The SR has been proposed as a model of how humans and other animals learn and generalize (Momennejad et al., , 2017; Stachenfeld et al., , 2017; Geerts et al., , 2020; Gershman, , 2018). For example, humans adapt more readily to changes in reward functions than changes in transition functions in a simple sequential binary decision-making task (Momennejad et al., , 2017), consistent with the SR facilitating rapid adaptation to changes in $\bm{r}$ but only slow learning of $\bm{M}$. Additionally, hippocampal ‘place cells’ have been proposed to encode a predictive map, with each cell corresponding to a column of $\bm{M}$ (Stachenfeld et al., , 2017). In this model, the firing of a place cell in a given location $s$ reflects the expected future occupancy of its ‘preferred’ location $s^{\prime}$ conditioned on currently being at $s$. Stachenfeld et al., (2017) showed that the SR model explains a range of findings in the hippocampal literature. For example, this model explains the asymmetry of place fields on a one-dimensional track (Mehta et al., , 2000), where $s^{\prime}$ is more likely to be reached from other states $s$ that precede $s^{\prime}$ than equidistant states that follow $s^{\prime}$. The SR also explains why place fields change near a newly inserted barrier, since two states on either side of the barrier can no longer be visited in quick succession (Alvernhe et al., , 2011).
While the SR is perhaps the most prominent model in systems neuroscience that combines features of model-free and model-based RL, it is not the only one. Another interesting algorithm is the ‘DYNA’ architecture of Sutton, (1991). In this framework, a model of the world is learned from experience and used to train a model-free policy offline by bootstrapping imagined experience sampled from the model. This allows for more data-efficient learning of model-free policies at the cost of additional compute during ‘rest’, but without needing more compute at decision time (Figure 4E). The model used to simulate data for offline training can either be an explicit learned world model, or it can simply be a memory buffer of past experiences in the form of $(s_{t},a_{t},r_{t},s_{t+1},a_{t+1})$ tuples. Such experience replay has also proven crucial to the success of modern deep reinforcement learning agents by allowing for higher data efficiency and reducing the instability arising from online experience being autocorrelated (Mnih et al., , 2013; Schaul et al., , 2015). A prominent theory in neuroscience posits that hippocampal replays could be implementing such a DYNA-like algorithm by generating imagined experience that is used to train the model-free RL systems of the brain (Mattar and Daw, , 2018). This theory is supported by the finding that patterns of rodent replay in multiple navigation tasks are consistent with the optimal replays of a Q-learning agent with DYNA, and it has recently been extended to explain not just the content of replay but also their timing (Agrawal et al., , 2022).

### v3

@@ 4 Q-learning
In some cases, we may not know the transition function or it could be expensive to simulate. Additionally, it has been found that dopamine activity can reflect learning signals not just for state values but also for action values [91, 75]. This suggests an alternative model of biological learning, where animals directly learn the state-action values defined in Equation 10. This is in contrast to the algorithm in Section 3, where the agent only learned the state values, and then computed the action values at decision-time using a one-step world model. Q-learning is the most prominent model for learning state-action values, and it has commonly been used to explain animal behaviour and neural activity [76, 64].
To learn the Q-values necessary for action selection directly, we start by expanding Equation 10,
@@
Importantly, this self-consistency expression only holds when the Q-values have converged to the true expected rewards, and the associated greedy policy is therefore optimal [108]. We can now use Equation 13 as an objective by defining
$\mathcal{L}\propto{\left(Q(s,a)-\left(r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim p(s^{\prime}|s,a)}\left[\text{max}_{a^{\prime}}Q(s^{\prime},a^{\prime})\right]\right)\right)}^{2},$ (14)
@@
This is the so-called Q-learning update rule (127; Figure 3A), where we have estimated the expectation with the single sample actually seen by the agent in the last line.
@@
Q-learning is guaranteed to converge to the optimal policy in the limit of infinitesimal learning rates and infinite sampling of the state-action space [126, 108]. However, following the greedy policy $a^{*}(s)=\text{argmax}_{a}Q(s,a)$ before convergence of the Q-values can lead to undersampling of the state-action space and poor performance. It is therefore common to either use an ‘$\epsilon$-greedy’ policy, $\pi(a|s)=\epsilon/|\mathcal{A}|+(1-\epsilon)I_{a=a^{*}(s)}$, or a softmax-policy, $\pi(a|s)\propto\text{exp}(\beta Q(s,a))$, to collect the experience used to update the Q-values (Figure 3B). Such exploration strategies and their biological correlates are discussed in more detail in Section 9.
These approaches make Q-learning an ‘off-policy’ algorithm, since the policy used in the learning update (the greedy policy) is different from the policy used for action selection (the stochastic policy). An on-policy alternative known as ‘SARSA’ (state-action-reward-state-action) is also commonly used, where the update rule uses the Q-value corresponding to the action $a_{t+1}$ sampled at the next timestep instead of the greedy action (Figure 3C):
@@
This will converge to the true Q-values for a given policy $\pi$, similar to how the TD learning rule in Equation 9 converges to the true value function for a given policy, again under assumptions of infinite sampling of the space.
When animals have to choose between actions with different values, studies have found evidence for midbrain dopamine neurons encoding the prediction error used for either Q-learning [91, 76] or SARSA [75, 76]. They may therefore not just be abstract learning algorithms, but instead have plausible implementations in biological neural circuits. However, the methods considered so far also have notable shortcomings. For example, the amount of data needed to learn state(-action) values and the assumption of a stationary environment can be prohibitive for animals needing to act in a rapidly changing world, where bad decisions have fatal consequences.
@@ 5 Model-free and model-based reinforcement learning
We have so far considered what is known as ‘model-free’ reinforcement learning algorithms. These involve learning a stimulus-response function that says ‘when in state $s$, take action $a$’. Such algorithms do not require much computation at decision time, where they rely on cached state or action values. However, it can require a lot of experience with the environment to learn these model-free policies, and they can be inflexible in changing environments. This is incompatible with many aspect of animal behaviour, which we know is adaptive and can benefit from ‘latent learning’ in an environment before a reward-driven task is ever encountered [12, 116].
On the other hand, ‘model-based’ reinforcement learning uses a model of the world to simulate the consequences of different actions at decision time. This can be much more data efficient, since learning a world model is often easier than learning a full policy (Figure 4A). In machine learning settings, model-based RL has exhibited impressive performance across a range of tasks with large state spaces, including Atari, chess, shogi, and Go [98, 95, 26]. In a biological context, the idea of first learning a model of the environment, and then using it to guide reward-driven behaviour also provides one plausible explanation for latent learning and other types of rapid adaptation. However, model-based decision making can be computationally intensive at decision-time, which is a challenge for animals that rely on rapid decision making for survival (Figure 4B).
@@
If the environment is determinstic, $p(s^{\prime}|s,a)$ is a delta function, and otherwise the next-state expectations may need to be approximated with multiple samples. Unfortunately, optimizing over all possible action sequences in Equation 18 is in general an exponentially large search problem in the planning depth, which makes it infeasible for any reasonably sized problem. It is therefore common to either use ‘depth-first search’ with limited breadth, or ‘breadth-first search’ with limited depth. In breadth-first search, we consider all possible actions at each level of the search tree but terminate the search at a finite depth, instead using cached ‘model-free’ state-values to estimate the reward-to-go from the terminal states. Such ‘plan-until-habit’ has also been proposed as a model of human behaviour [52]. In depth-first search, we instead sample a series of paths from $s_{t}$ to termination (or some upper bound), using a heuristic to prioritize actions expected to be good, and then pick an action with high expected reward [45].
@@
For both of these strategies, it is necessary to trade off the temporal opportunity cost of planning with the increase in expected reward [16, 1]. This has been a popular research area in cognitive science, where a wealth of literature on ‘resource-rational’ decision making has emerged in recent years [38, 18]. However, this literature has often focused on the behaviour of optimal agents, with less focus on the learning process and neural mechanisms that might implement the necessary computations. Bridging this gap, recent work has suggested that frontal cortex and striatum might initially store a ‘model-free’ policy in its network state, which is gradually updated with model-based information from the hippocampal formation until the policy improvement is outweighed by the temporal opportunity cost of planning [47].
@@
While several model-based and model-free reinforcement learning methods have thus been developed and used to model animal learning and behaviour, it remains an open question when and whether these different strategies drive animal behaviour. A different line of research has therefore explicitly investigated the balance between model-based and model-free RL in biological agents [23, 35, 55], where the choice between the two approaches is thought to be guided by some notion of optimality on the basis of available resources and uncertainty about the environment. A popular paradigm for these studies is the so-called ‘two-step’ task developed by Daw and colleagues (22, 72, 123; although note 3).
@@
Such work has shown that animals can use both model-free and model-based decision making, with the dorsolateral striatum being particularly important for model-free reinforcement learning [132, 133], and the dorsomedial striatum, prefrontal cortex, and hippocampal formation being important for model-based decision making [120, 35, 68, 76, 53]. This also has interesting parallels to recent work in motor learning, where the basal ganglia were found to be sufficient for ‘habitual’ motor sequences even in the absence of motor cortex, while motor cortex was necessary for more flexible motor behaviours that are likely to require a high-level ‘schema’ of the task structure [70, 69]. In Section 9 we will see how combining these model-based and model-free ideas with deep learning can lead to human-level performance in tasks such as chess and Go that require long-term planning.
@@ 6 The successor representation
As we saw in the previous section, an important distinction can be made between model-free reinforcement learning methods, which cache stimulus-response mappings based on prior experience, and model-based reinforcement learning methods, which compute a policy by simulating possible futures using a world model at decision-time. However, we have also noted how animals both need the flexibility of model-based methods as well as the rapid decision making afforded by model-free methods. It has therefore been suggested that animals use intermediate methods that combine some model-free and some model-based features. A particularly prevalent theory has been that of the ‘successor representation’ (SR), which has been proposed to explain both human behaviour [72] and features of neural activity [103]. In particular, the SR allows for flexible adaptation to changing reward functions without having to perform expensive simulations at decision time.
@@
Figure 4: Model-based reinforcement learning. (A) Learning curves for model-free (MF) and model-based (MB) RL agents. The MB agent used depth-first search to compute an optimal path at each decision point, gradually learning the reward and transition functions while exploring the environment. The MF agent was a Q-learning agent with $\epsilon=0$ and learning rate $\alpha=1$. (B) Wallclock time needed to run 100 episodes of cliffworld with either the MB or MF agents from (A), as a function of the length of the environment. While the MB agent required less experience to learn a good policy, the wallclock time per episode was much larger than for the MF agent. This illustrates an important balance between model-based and model-free reinforcement learning, where MF methods usually require more experience but MB methods require more compute at decision time. (C) Learning curve for an agent using the successor representation (SR) together with learning curves for the model-based agent in (A) and the greedy TD-agent from Figure 2. The goal was moved from location (9, 0) to location (0, 4) at episode 40 (vertical black line), and location (9, 0) was instead given a reward of -5. The MB and SR agents had their reward functions updated to reflect this change and rapidly adapted their policies, while the TD agent had no such mechanism for robustness to changing reward functions. Reward curves were convolved with a Gaussian kernel ($\sigma=3$ episodes), which is why performance appears to decrease slightly before episode 40. The TD and SR agents were assumed to have access to a 1-step world model at initialization, while the MB agent learned the transition structure from experience. (D) SR agents cannot always adapt to new reward functions if the newly rewarded states have low probablity under the old policy. Left column: Value function for an agent that learned an initial policy in an environment with a small reward in the upper left corner and intermediate reward in the upper right corner. The middle top and bottom states are ‘cliffs’. The agent learned to make an initial rightward choice (grey arrows). Right column: A large reward was introduced in either the top left corner (top row) or bottom left corner (bottom row) and the value function recomputed (Equation 27). The agent was unable to adapt to a large reward in the bottom left corner, since the old policy had low probability of reaching this state, even after initially going to the left. This results in a low expected value for going left from the start state (red circle), and a suboptimal policy that continues to go right (red arrow). (E) Learning curve for a standard Q-learning agent (blue) or Dyna agents that perform different numbers of Q-value updates after each physical action (legend). These Dyna updates used cached experience rather than data from a learned world model. Dyna agents make better use of limited experience at the cost of increased compute (proportional to the number of updates).
The SR [24] rewrites the expected reward starting from state $s$ as:
@@
This provides a better starting point than the old policy and reward function (Figure 4C), but the SR does not generalize perfectly since the new value function can lead to policy changes and therefore having to update $\bm{M}$ (Figure 4D). The successor matrix can be learned by temporal-difference learning when transitioning from $s_{t}$, analogous to Equation 9:
@@
The SR has been proposed as a model of how humans and other animals learn and generalize [72, 103, 35, 36]. For example, humans adapt more readily to changes in reward functions than changes in transition functions in a simple sequential binary decision-making task [72], consistent with the SR facilitating rapid adaptation to changes in $\bm{r}$ but only slow learning of $\bm{M}$. Additionally, hippocampal ‘place cells’ have been proposed to encode a predictive map, with each cell corresponding to a column of $\bm{M}$ [103]. In this model, the firing of a place cell in a given location $s$ reflects the expected future occupancy of its ‘preferred’ location $s^{\prime}$ conditioned on currently being at $s$. Stachenfeld et al. [103] showed that the SR model explains a range of findings in the hippocampal literature. For example, this model explains the asymmetry of directional place fields on a one-dimensional track [65], where $s^{\prime}$ is more likely to be reached from other states $s$ that precede $s^{\prime}$ than equidistant states that follow $s^{\prime}$. The SR also explains why place fields change near a newly inserted barrier, since two states on either side of the barrier can no longer be visited in quick succession [5].
While the SR is perhaps the most prominent model in systems neuroscience that combines features of model-free and model-based RL, it is not the only one. Another interesting algorithm is the ‘Dyna’ architecture of Sutton [106]. In this framework, a model of the world is learned from experience and used to train a model-free policy offline by bootstrapping imagined experience sampled from the model. This allows for more data-efficient learning of model-free policies at the cost of additional compute during ‘rest’, but without needing more compute at decision time (Figure 4E). The model used to simulate data for offline training can either be an explicit learned world model, or it can simply be a memory buffer of past experiences in the form of $(s_{t},a_{t},r_{t},s_{t+1},a_{t+1})$ tuples. Such experience replay has also proven crucial to the success of modern deep reinforcement learning agents by allowing for higher data efficiency and reducing the instability arising from online experience being autocorrelated [71, 94]. A prominent theory in neuroscience posits that hippocampal replays could be implementing such a Dyna-like algorithm by generating imagined experience that is used to train the model-free RL systems of the brain [64]. This theory is supported by the finding that patterns of rodent replay in multiple navigation tasks are consistent with the optimal replays of a Q-learning agent with Dyna, and it has recently been extended to explain not just the content of replays but also their timing [1].